# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v2

@@ IV-B Policy Optimization
$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
Robot Navi. $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
TABLE II: Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics. The best results are bolded.
Conciseness Consistency Strictness Accuracy
STONE STONE STONE STONE
Scenarios Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5
CtF Capture 0.41 0.21 0.20 0.22 0.33 0.08 0.08 0.13 0.33 0.08 0.08 0.13 0.95 0.95 0.95 0.94
CtF Fight 0.39 0.18 0.25 0.25 0.68 0.09 0.50 0.50 0.68 0.09 0.50 0.50 0.94 0.92 0.94 0.93
CtF Patrol 0.38 0.18 0.25 0.28 0.63 0.05 0.50 0.53 0.63 0.05 0.50 0.53 0.95 0.93 0.94 0.94
CtF Mixture 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
CtF Capt. 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
Fetch Push 0.42 0.31 0.25 0.25 1.00 0.28 0.50 0.50 1.00 0.28 0.50 0.50 0.99 0.97 0.97 0.97
Robot Navi. 0.48 0.22 0.25 0.25 0.68 0.15 0.50 0.50 0.68 0.15 0.50 0.50 0.97 0.96 0.96 0.96
@@ IV-C Experimental Results
IV-C1 Benefits and Limitations of TLNet
Table I presents the most frequent explanation obtained by each method for each scenario. Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively easy-to-understand explanations. In comparison, NN-TLI typically yielded verbose explanations that are difficult to understand, while STONE often captured only one temporal clause, resulting in incomplete explanations.
The explanations produced by TLNet are also qualitatively reasonable (and often correct). For example, for all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario. The inferred CtF constraints are also syntactically reasonable. For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which required the blue agent to learn to prevent the red agent from reaching the blue flag. Other scenarios, such as Fight and Patrol, did not include this predicate in the inferred constraint because those red policies do not target the blue flag. In comparison, the CtF constraints inferred by NN-TLI are generally too verbose to understand, while STONE did not infer any constraints.
For Fetch Push, TLNet also correctly inferred the task clause as $\psi_{\mathrm{bt}}$ (“block reaches the target”), which no other method correctly inferred. While TLNet inferred part of the expected constraint ($\psi_{\mathrm{gb}},\psi_{\mathrm{bt}}$), it missed an expected constraint $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”), likely due to insufficient occurrences of the grip touching the obstacle in the sampled data. Future research could mitigate this issue by introducing more intelligent sampling strategies for generating the negative dataset (e.g., adversarial or targeted trajectory sampling). NN-TLI also inferred part of the expected constraint, but also missed the $\neg\psi_{\mathrm{go}}$ constraint. Meanwhile, both variations of STONE inferred an incorrect constraint of $\psi_{\mathrm{bt}}$ (“block reaches the target”), since this is actually the task and not possible to always satisfy.
@@ IV-C1 Benefits and Limitations of TLNet
As shown in Table II, TLNet also outperformed all baselines in conciseness, consistency, and strictness, with marginal differences in classification accuracy. Across all scenarios, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness. These results support the qualitative findings above, validating the explainability of our method’s outputs and the use of these explainability metrics.
@@ IV-C2 Benefit of Including Weights
We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [25] or LTL [16, 17, 18]) by examining how inferred weights differ across CtF scenarios. Consider the Capture and Mixture explanations in Table I. Both capture and mixture red policies target the blue flag, but the capture policy is more aggressive in doing so. The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mixture, focusing more on capturing the red flag. This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mixture explanation (0.56), due to the Capture policy’s prioritization of flag protection. Conversely, the Mixture policy shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture policy (0.30), due to the Mixture policy’s prioritization of flag capture. Note that $\psi_{\mathrm{ba,rf}}$ appears in both the task and constraint clauses in these explanations. This duplication results from our design choice to fix equal weights (0.5) on the task and constraint clauses, which forces TLNet to express prioritization between the task and constraint by including relevant task predicates in the constraint clause. Although one could alternatively include the task and constraints weights within the optimized parameters, we found that doing so introduced training instability.
@@
TABLE III: Ablation Study of TLNet in the CtF Capture Scenario.
Metrics Full No Filter No Regularization No Pruning
Conciseness 0.408 0.289 0.396 0.147
Consistency 0.325 0.175 0.233 0.2
Strictness 0.117 0.106 0.117 0.063
Accuracy 0.951 0.947 0.954 0.776
@@ IV-C3 Importance of TLNet Components
Finally, the ablation study results shown in Table III show that filtering, regularization, and pruning are all important to the performance of TLNet. Removing predicate filtering notably reduced conciseness (by 27.2%) and consistency (by 46.1%). Removing regularization decreased conciseness (by 2.94%) and significanbly reduced consistency (by 28.3%) with marginal accuracy improvement. Finally, removing the pruning step nearly tripled conciseness and halved strictness.
@@ V Conclusion
We introduce a neuro-symbolic framework, based on the TLNet architecture, for generating wSTL explanations of robot policies. Our approach generates easy-to-understand explanations through predicate filtering, loss regularization, and weight pruning. To assess explanation quality beyond classification metrics, we also propose explainability metrics—conciseness, consistency, and strictness. Experiments in three simulated robotic environments show that our method produces explanations that better capture these metrics while retaining high classification accuracy. By bridging neural network learning and formal logic-based explanations, our work advances explainability in complex robotic applications, with future directions including more expressive wSTL structures, human-in-the-loop refinement, and explainability-by-design policy optimization.

### v3

@@ IV-B Policy Optimization
40%
$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
100%
$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
100% $\mathcal{G}_{I}[\neg\psi_{\mathrm{og}}]$ 100%
Robot
Nav.
$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge$
$0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$
70% $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ 70%
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
100%
$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
100%
$\mathcal{F}_{I}[\psi_{\mathrm{eg}}\vee$
$\psi_{\mathrm{ec}}\vee\psi_{\mathrm{eh}}]$
100%
TABLE III: Comparison of Conciseness, Consistency, Strictness with standard deviation. The best results are bolded. [^5] Statistical significance (Friedman test, Dunn-Bonferroni analysis) against the second best values indicated with * ($p<0.05$), ** ($p<0.01$).
Conciseness Consistency Strictness
STONE STONE STONE
Scenario Ours NN-TLI top-3 top-5 DT Ours NN-TLI top-3 top-5 DT Ours NN-TLI top-3 top-5 DT
[…]
TABLE IV: Comparison of Accuracy. The best results are bolded.
Accuracy
STONE
Scenario Ours NN-TLI top-3 top-5 DT
CtF Capt. $\textbf{0.95}_{\pm 0.02}$ $\textbf{0.95}_{\pm 0.01}$ $\textbf{0.95}_{\pm 0.01}$ $0.94_{\pm 0.02}$ 0.95±0.00
CtF Fight 0.94±0.02 0.92±0.02 0.94±0.02 0.93±0.01 0.94±0.00
CtF Patrol 0.95±0.02 0.93±0.02 0.94±0.01 0.94±0.02 0.95±0.00
CtF Mix 0.87±0.01 0.89±0.04 0.88±0.03 0.88±0.05 0.89±0.00
CtF Capt.0 0.95±0.02 0.95±0.01 0.95±0.01 0.95±0.02 0.95±0.00
[…]
Robot Nav. 0.97±0.01 0.96±0.01 0.96±0.01 0.96±0.01 0.98±0.01
@@ IV-C Experimental Results
IV-C1 Benefits & Limitations of TLNet
Table II presents the most frequent explanation obtained by each method for each scenario. Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively comprehensible explanations. In comparison, NN-TLI typically yielded verbose explanations that are difficult to understand, while STONE often captured only one temporal clause, resulting in incomplete explanations.
The explanations produced by TLNet are also qualitatively reasonable. For all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario. The inferred CtF constraints are also syntactically reasonable. For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which requires the blue agent to learn to prevent the red agent from reaching the blue flag. Other scenarios, such as Fight and Patrol, did not include this predicate in the inferred constraint because those red policies do not target the blue flag. In comparison, the CtF constraints inferred by NN-TLI are generally too verbose to understand, while STONE did not infer any constraints. The DT (STL) baseline often produced a task- or constraint-only STL explanation, typically with semantically-inaccurate clauses. For Fetch Push, TLNet also correctly inferred the task clause as $\psi_{\mathrm{bt}}$ (“block reaches the target”), which no other method correctly inferred. While TLNet inferred part of the expected constraint ($\psi_{\mathrm{gb}},\psi_{\mathrm{bt}}$), it missed an expected constraint $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”), likely due to insufficient occurrences of the grip touching the obstacle in the sampled data. NN-TLI also inferred part of the expected constraint, but also missed the $\neg\psi_{\mathrm{go}}$ constraint. Meanwhile, both variations of STONE inferred an incorrect constraint of $\psi_{\mathrm{bt}}$ (“block reaches the target”), since this is actually the task and not possible to always satisfy.
@@ IV-C1 Benefits & Limitations of TLNet
[…]
@@ IV-C2 Benefit of Including Weights
We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [15, 6, 1] or LTL [10]) by examining how inferred weights differ across CtF scenarios. Consider the Capture and Mix explanations in Table II. Both capture and mix red policies target the blue flag, but the capture policy is more aggressive in doing so. The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mix, focusing more on capturing the red flag. This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mix explanation (0.56), due to the Capture policy’s prioritization of flag protection. Conversely, the Mix explanation shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture explanation (0.30), due to its prioritization of flag capture. Note that $\psi_{\mathrm{ba,rf}}$ appears in both the task and constraint clauses in these explanations. This duplication results from our design choice to fix equal weights (0.5) on the task and constraint clauses, which forces TLNet to express prioritization between the task and constraint by including relevant task predicates in the constraint clause.
@@
TABLE V: Ablation Study of TLNet in the CtF Capture Scenario.
Features Concise. Consist. Strict. Accuracy
[…]
No Pruning: $_{N_{\mathrm{w}}=0}$ 0.15±0.02 0.20 0.06±0.00 0.78±0.05
[…]
@@ IV-C3 Importance of TLNet Components
Finally, the ablation study results shown in Table V show that filtering, regularization, and pruning are all important to the performance of TLNet. Removing predicate filtering notably reduced conciseness and consistency, removing regularization decreased conciseness and significantly reduced consistency with no accuracy improvement, and removing pruning largely reduced all metrics. We also tested an argmin approach to sample negative trajectories, where rollouts used the lowest-probability action in the target policy, and saw no significant difference in results compared to random sampling.
@@ V Conclusion
We introduce TLNet, a neuro-symbolic framework for generating wSTL explanations of robot policies through predicate filtering, loss regularization, and weight pruning. We also propose novel explainability metrics—conciseness, consistency, and strictness—to assess explanation quality beyond classification accuracy. Experiments in three simulated robotic environments show our method outperforms baselines on these metrics while retaining high classification accuracy. Future directions include real-world deployment with online inference (which would require improving sample efficiency and handling predicate uncertainty from sensor noise and partial observability), learning task-constraint weights with a stabilizer, improving negative data generation, exploring post-hoc enforcement of the task-constraint format, and user studies.